# Identification and Scaling of Latent Variables in Ordinal Factor Analysis

### v2

@@ Abstract
This work was supported by the Institute of Education Sciences, U.S. Department of Education, Grant R305D210044. Correspondence to Edgar Merkle, 219 McAlester Hall, Columbia, Missouri, USA. Email: merklee@missouri.edu
@@
Researchers have long considered problems with the treatment of ordinal variables as continuous [<]e.g.,>bolbar81,burvuo19,lidkru18,mcn20,rhebro12,winmar84. Perhaps the most famous work on this topic is Stevens’ scales of measurement [<]e.g.,>ste46. In distinguishing between ordinal scales and interval scales, Stevens notes that “means and standard deviations computed on an ordinal scale are in error to the extent that the successive intervals on the scale are unequal in size” (p. 679). Our results below involve the idea of equal intervals in ordinal CFA models, providing minimal identification constraints that are related to equal intervals. Our results are also related to those of Kruschke and colleagues [kru14, kru15, lidkru18], who considered identification constraints for univariate, ordinal regression models. They reasoned that, because applied researchers are accustomed to treating ordinal variables as continuous, we should seek to identify the ordinal regression model so that the underlying continuous variable is on the scale of the ordinal variable. For example, if we have an ordinal variable with 5 categories, then the ordinal regression model should generally predict values between 1 and 5 on the latent continuous scale, which are then converted to probabilities of assuming each ordered category.
@@ 1.2 Latent Variable Prediction
Following model estimation via marginal maximum likelihood or weighted least squares, researchers may optionally request latent variable predictions that serve as scores for each individual $i$. These predictions can be obtained by maximizing the likelihood function $L(\bm{\eta}_{i}\mid\bm{y}_{i},\bm{\xi})$ for all $i$, where the likelihood function is defined by the right side of Equation (5). As compared to Equation (5), we now estimate $\bm{\eta}_{i}$ and condition on $\bm{\xi}$ whereas we previously did the opposite. Maximization of this function requires numerical methods because it involves the normal CDF.
@@ 2.2 Empirical Results
As described previously, the latent variable predictions involve maximization of Equation (5), which is now a function of $\bm{\eta}_{i}$ and is conditioned on $\bm{\xi}$ along with $\bm{y}_{i}$. To show that the model constraints from the previous section lead to latent variable predictions equaling the average of observed variables, we consider here a one-factor model with values of $p$ from 2 to 10 and $K=5$. For each value of $p$, we generate all possible response patterns and calculate the empirical Bayes prediction of the latent variable for each response pattern. We do not estimate any item parameters here: in addition to the constraints on the loadings, thresholds, and latent means from the previous subsection, we fixed $\bm{\nu}=\bm{0}$, $\bm{\Phi}=p$, and $\bm{\Theta}=\bm{I}$.
@@
Figure 1 shows scatter plots of the average observed response (x-axis) versus empirical Bayes latent variable prediction for all possible response patterns. Each red point is a response pattern that does not include an extreme response of 1 or 5, while each blue point is a response pattern that does include an extreme. The figure shows that the points generally fall along the diagonal, with some differences at the far left and far right side of each panel. This provides some evidence that latent variable predictions under our model constraints remain close to the average of observed variables for all response patterns. The supplementary materials include additional code that considers additional values of $p$ and $K$. It shows that the gradient of the likelihood function is always close to 0 at the mean of observed variables, so long as the response pattern does not include extreme responses of 1 or $K$. The code also considers maximum likelihood estimates of the latent variables, in addition to empirical Bayes predictions.
@@
When an individual’s response pattern does include the extremes of 1 or $K$, the $\bm{\eta}_{i}$ predictions are pulled toward -$\infty$ or $\infty$, respectively, so that they no longer equal the average of the observed variables. This can be observed on the left and right sides of each panel of Figure 1. A similar phenomenon happens for large values of $K$ (say 8 or more) when responses are near the extremes (e.g., 2 or (K1)(K-1)). In the empirical Bayes case, the prior distribution from (2) helps keep the predictions from straying too far from the observed average. We fixed the prior variance, $\bm{\Phi}$, to equal $p$ in each panel of these results. This may appear to be an odd choice, but it is also a weaker prior variance as compared to a value of 1. Were we to fix $\bm{\Phi}$ to 1, our points would be even closer to the diagonal. And in the maximum likelihood case, an adhoc, vague prior distribution is often used to ensure that latent variable predictions exist for extreme response patterns. In either case, the resulting latent variable predictions are close to the means of the integer-coded ordinal variables.
@@
Figure 1: Observed averages versus empirical Bayes latent variable estimates for $K=5$ and $p=2$ to 10. Each point represents a response pattern. Red points are response patterns that do not include a response of 1 or 5, and blue points are response patterns that do include a response of 1 and/or 5.
@@ 5.2 Results
We did not find much evidence that the integer constraints had estimation differences as compared to alternative identification constraint methods. Minor differences in convergence rates existed, but these were balanced out by differences in admissible result rates, resulting in almost identical converged and admissible (i.e., valid) result rates. Results of an ANOVA with converged and admissible result rates as the outcome variable indicated that the identification constraint had a negligible effect ($\eta^{2}_{p}=0.001$). Similarly, starting values also minimally affected converged and admissible result rates ($\eta^{2}_{p}=0$). Other simulation factors had a larger impact, ranging from $\eta^{2}_{p}=0.043$ for response distribution to $\eta^{2}_{p}=0.229$ for factor loading magnitude. Given the minimal impact of starting values, we will focus on the results when using simple starting values. Results for default starting values are presented in Appendix B.
@@ 6.1 Method
After model estimation, we obtained empirical Bayes estimates of the latent variable for each respondent.
@@
Figure 3: Average of observed variables versus empirical Bayes latent variable predictions for the attitudes toward science dataset.
@@ 6.2 Results
Finally, Figure 3 compares the average of each participant’s ordinal variables to the empirical Bayes prediction of the latent variable under the alternative constraints. We see that the empirical Bayes predictions are similar to the averages, with some shrinkage whereby the extreme averages have less-extreme latent variable predictions. We also see that the averages and latent variable predictions differ the most for participants with low averages (below 2), reflecting the result that participants generally tended to respond with “agree” or “strongly agree” on the ordinal scale.
@@ 7.2 Results
Figure 4: Average of observed variables versus empirical Bayes latent variable predictions for the social media dataset.
Figure 4 is similar to Figure 3 from our previous example, showing the average score for each individual versus the empirical Bayes predictions from the integer-constrained model. We see close agreement here, with points falling slightly below the diagonal for average scores near 3 and larger. This is related to our observation that the thresholds for all items are below the benchmarks of 2.5 and 3.5: the model estimates that people tend to avoid the “strongly disagree” option, so lower values of the latent variable can still lead participants to select higher response options.
Figure 5: Latent variable values versus expected average score (left panel), with overlaid points of empirical Bayes latent variable estimates versus observed average scores (right panel).
@@ 8.3 Summary
All results were obtained using the R system for statistical computing [r23], version 4.5.1, making use of the lavaan [ros12], ggplot2 [ggplot2], and xtable [xtable] packages. R is freely available under the General Public License 2 from the Comprehensive R Archive Network at http://CRAN.R-project.org/.
@@ Appendix A Alternative Integer Constraints
Under this constraint, we experienced inconsistent signs of individual loadings. For example, when we have an even number of loadings, the loadings can all be negative yet have a geometric mean of 1. So researchers may experience problems with sign indeterminacy under these constraints.

### v3

@@ Abstract
Correspondence to Edgar Merkle, 219 McAlester Hall, Columbia, Missouri, USA. Email: merklee@missouri.edu
Financial support: This work was supported by the Institute of Education Sciences, U.S. Department of Education, Grant R305D210044.
Competing interests: The authors declare none.
@@
Many researchers have studied when and whether we can treat ordinal variables as continuous [<]e.g.,>bolbar81,burvuo19,lidkru18,mcn20,rhebro12,winmar84. Perhaps the most famous work on this topic is Stevens’ scales of measurement [<]e.g.,>ste46. In distinguishing between ordinal scales and interval scales, Stevens notes that “means and standard deviations computed on an ordinal scale are in error to the extent that the successive intervals on the scale are unequal in size” (p. 679). Our results below involve the idea of equal intervals in ordinal CFA models, providing minimal identification constraints that are related to equal intervals. Our results are also related to those of Kruschke and colleagues [kru14, kru15, lidkru18], who considered identification constraints for univariate, ordinal regression models. They reasoned that, because applied researchers are accustomed to treating ordinal variables as continuous, we should seek to identify the ordinal regression model so that the underlying continuous variable is on the scale of the ordinal variable. For example, if we have an ordinal variable with 5 categories, then the ordinal regression model should generally predict values between 1 and 5 on the latent continuous scale, which are then converted to probabilities of assuming each ordered category.
@@ 1.2 Latent Variable Prediction
[…]
For ordinal factor analysis, we can obtain latent variable predictions by maximizing the likelihood function $L(\bm{\eta}_{i}\mid\bm{y}_{i},\bm{\xi})$ for all $i$, where the likelihood function has the same form as the right side of Equation (5). As compared to Equation (5), we now estimate $\bm{\eta}_{i}$ and condition on $\bm{\xi}$ whereas we previously did the opposite. Maximization of this function requires numerical methods because it involves the normal CDF.
@@ 2.2 Empirical Results
As described previously, the latent variable predictions involve maximization of Equation (5), which is now a function of $\bm{\eta}_{i}$ and is conditioned on $\bm{\xi}$ along with $\bm{y}_{i}$. To show that the model constraints from the previous section lead to latent variable predictions equaling the average of observed variables, we consider here a one-factor model with values of $p$ from 2 to 10 and $K=5$. For each value of $p$, we generate all possible response patterns and calculate the MAP prediction of the latent variable for each response pattern. We do not estimate any item parameters here: in addition to the constraints on the loadings, thresholds, and latent means from the previous subsection, we fixed $\bm{\nu}=\bm{0}$, $\bm{\Phi}=p$, and $\bm{\Theta}=\bm{I}$.
@@
Figure 1 shows scatter plots of the average observed response (x-axis) versus MAP latent variable prediction for all possible response patterns. Each red point is a response pattern that does not include an extreme response of 1 or 5, while each blue point is a response pattern that does include an extreme. The figure shows that the points generally fall along the diagonal, with some differences at the far left and far right side of each panel. This provides some evidence that latent variable predictions under our model constraints remain close to the average of observed variables for all response patterns. The supplementary materials include additional code that considers additional values of $p$ and $K$. It shows that the gradient of the likelihood function is always close to 0 at the mean of observed variables, so long as the response pattern does not include extreme responses of 1 or $K$. The code also considers maximum likelihood estimates of the latent variables, in addition to MAP predictions.
@@
When an individual’s response pattern does include the extremes of 1 or $K$, the $\bm{\eta}_{i}$ predictions are pulled toward -$\infty$ or $\infty$, respectively, so that they no longer equal the average of the observed variables. This can be observed on the left and right sides of each panel of Figure 1. A similar phenomenon happens for large values of $K$ (say 8 or more) when responses are near the extremes (e.g., 2 or (K1)(K-1)). In the MAP case, the prior distribution from (2) helps keep the predictions from straying too far from the observed average. We fixed the prior variance, $\bm{\Phi}$, to equal $p$ in each panel of these results. This may appear to be an odd choice, but it is used here to demonstrate robustness of our result. This is because a prior variance of $p$ is weaker than the traditional prior variance of 1. Were we to fix $\bm{\Phi}$ to 1, our points would be even closer to the diagonal. And in the maximum likelihood case, an adhoc, vague prior distribution is often used to ensure that latent variable predictions exist for extreme response patterns. In each case, the resulting latent variable predictions are close to the means of the integer-coded ordinal variables.
@@
Figure 1: Observed averages versus MAP latent variable estimates for $K=5$ and $p=2$ to 10. Each point represents a response pattern. Red points are response patterns that do not include a response of 1 or 5, and blue points are response patterns that do include a response of 1 and/or 5.
@@ 5.2 Results
We did not find much evidence that the integer constraints had estimation differences as compared to alternative identification constraint methods. Minor differences in convergence rates existed, but these were balanced out by differences in admissible result rates, resulting in almost identical converged and admissible (i.e., valid) result rates. Results of an ANOVA with converged and admissible result rates as the outcome variable indicated that the identification constraint had a negligible effect ($\eta^{2}_{p}=0.001$). Similarly, starting values also minimally affected converged and admissible result rates ($\eta^{2}_{p}=0$). Other simulation factors had a larger impact, ranging from $\eta^{2}_{p}=0.088$ for response distribution to $\eta^{2}_{p}=0.238$ for factor loading magnitude. Given the minimal impact of starting values, we will focus on the results when using simple starting values. Results for default starting values are presented in Appendix B.
@@ 6.1 Method
After model estimation, we obtained MAP estimates of the latent variable for each respondent.
@@
Figure 3: Average of observed variables versus MAP latent variable predictions for the attitudes toward science dataset.
@@ 6.2 Results
Finally, Figure 3 compares the average of each participant’s ordinal variables to the MAP prediction of the latent variable under the alternative constraints. We see that the MAP predictions are similar to the averages, with some shrinkage whereby the extreme averages have less-extreme latent variable predictions. We also see that the averages and latent variable predictions differ the most for participants with low averages (below 2), reflecting the result that participants generally tended to respond with “agree” or “strongly agree” on the ordinal scale.
@@ 7.2 Results
Figure 4: Average of observed variables versus MAP latent variable predictions for the social media dataset.
Figure 4 is similar to Figure 3 from our previous example, showing the average score for each individual versus the MAP predictions from the integer-constrained model. We see close agreement here, with points falling slightly below the diagonal for average scores near 3 and larger. This is related to our observation that the thresholds for all items are below the benchmarks of 2.5 and 3.5: the model estimates that people tend to avoid the “strongly disagree” option, so lower values of the latent variable can still lead participants to select higher response options.
Figure 5: Latent variable values versus expected average score (left panel), with overlaid points of MAP latent variable estimates versus observed average scores (right panel).
@@ 8.3 Summary
All results were obtained using the R system for statistical computing [r23], version 4.5.2, making use of the lavaan [ros12], ggplot2 [ggplot2], and xtable [xtable] packages. Code for reproducing the results in this paper and for applying integer constraints to other lavaan models is available at https://semtools.r-forge.r-project.org/.
@@ Appendix A Alternative Integer Constraints
Under this constraint, we may experience inconsistent signs of individual loadings. For example, when we have an even number of loadings, the loadings can all be negative yet have a geometric mean of 1. This will not necessarily be problematic in situations where all loadings are expected to have the same sign, and where we use frequentist estimation. It may be more problematic for Bayesian models where some loadings’ posterior distributions overlap with 0. In that case, MCMC methods may experience problems because they will sometimes encounter loadings with differing signs.